### v2

# Long-term foehn reconstruction combining unsupervised and supervised learning Reto Stauffer, Achim Zeileis, Georg J. Mayr

Long-term foehn reconstruction combining unsupervised and supervised learning
Reto Stauffer, Achim Zeileis, Georg J. Mayr
[…]
Foehn winds, characterized by abrupt temperature increases and wind speed
changes, significantly impact regions on the leeward side of mountain
ranges, e.g., by spreading wildfires. Understanding how foehn occurrences
change under climate change is crucial. Unfortunately, foehn cannot be
measured directly but has to be inferred from meteorological measurements
employing suitable classification schemes. Hence, this approach is typically
limited to specific periods for which the necessary data are available.
We present a novel approach for reconstructing historical foehn occurrences
using a combination of unsupervised and supervised probabilistic statistical
learning methods. We utilize in-situ measurements (available for recent
decades) to train an unsupervised learner (finite mixture model) for
automatic foehn classification. These labeled data are then linked to
reanalysis data (covering longer periods) using a supervised learner (lasso
or boosting). This allows to reconstruct past foehn probabilities based
solely on reanalysis data.
Applying this method to ERA5 reanalysis data for six stations across
Switzerland and Austria achieves accurate hourly reconstructions of north
and south foehn occurrence, respectively, dating back to 1940. This paves the way
for investigating how seasonal foehn patterns have evolved over the past
83 years, providing valuable insights into climate change impacts on these
critical wind events.
[…]
Reto Stauffer
Universität Innsbruck
Faculty of Economics and Statistics
and Digital Science Center
Universitätsstr. 15
6020 Innsbruck, Austria
E-mail:
1 Introduction
Foehn winds are downslope winds on the leeward side of mountains and can be found all around the world in areas with pronounced topographical features that impede the airflow, such as the European Alps, the Southern Alps in New Zealand, mountain ranges along the Mediterranean Sea, or the Rocky Mountains. Depending on the region, these winds are given specific names such as Santa Anna winds (Southern California), Chinook (Rocky Mountains), Mistral (Southern France), Bora (Croatia), or Foehn (Central Europe and New Zealand). For historical reasons ‘foehn’ has become a synonym for this type of terrain-induced wind phenomena.
Often, foehn is characterized by a sharp increase in wind speed and sudden changes in temperature and relative humidity, which can have a strong influence on the local climate and the people living in the affected areas. While foehn is often associated with a mild (and typically dry) climate, strong foehn events can also cause extensive damage to vegetation and man-made structures. In some areas it is not uncommon for strong foehn gusts to overturn trucks or vans, or for airports and harbors to be closed due to unsafe conditions. In addition, the strong and dry winds can kindle and spread domestic fires and wild fires (Schoennagel et al., 2004; Reinhard et al., 2005; Zumbrunnen et al., 2009) or affect the development of Antarctic ice shelves (e.g., Cape et al. 2015; Elvidge et al. 2020).
Although the conceptual model of foehn is well established (Armi and Mayr, 2007; Mayr and Armi, 2008; Armi and Mayr, 2011; Richner and Hächler, 2013), it cannot be measured directly. During the last decades several (semi)-automatic methods have been developed which allow to differentiate ‘foehn’ and ‘no-foehn’ events based on in-situ measurements from automated weather stations (AWSs). Among the frequently used algorithms are Widmer’s föhn index (Widmer, 1966; Courvoisier and Gutermann, 1971) based on Fisher’s linear discriminant analysis to distinguish between two or more distinct classes, and enhanced versions of it (e.g., Jansing et al. 2022). Other studies use decision-based or tree-based methods to classify foehn events (e.g., Dürr 2008; Speirs et al. 2013; Cape et al. 2015; Turton et al. 2018; Datta et al. 2019; Elvidge et al. 2020; Laffin et al. 2021; Francis et al. 2023; Laffin et al. 2021; Francis et al. 2023), all of which are deterministic methods, where the thresholds have often been selected manually. To overcome these limitations, Plavcan et al. (2014) proposed a method based on finite mixture models for automatic and fully parametric probabilistic foehn classification. To perform the classification, all methods require AWS measurements with high temporal resolution (ideally sub-hourly), which are typically only available for recent decades. While this allows to classify and analyze foehn when the AWS provides sufficient data, there is no information on foehn prior to the installation of the AWS, nor when there have been outages or the AWS has been decommissioned.
Additional information on the atmospheric conditions from numerical (re-)analysis models are an excellent source to complement the in-situ measurements. Reanalysis data sets are typically produced by physically based numerical weather prediction (NWP) models and sophisticated data assimilation schemes that use all available observations to estimate the ‘best known’ atmospheric state. An example is the global reanalysis data set ERA5 (Hersbach et al., 2023a, b) from the European Centre for Medium-Range Weather Forecasts (ECMWF). ERA5 provides global hourly 4-dimensional atmospheric conditions back to 1940. However, this comes with its own challenges. First, foehn is not explicitly modeled by the NWP models. Second, due to technical and computational limitations, the reanalysis can only approximate the real world and cannot resolve small-scale atmospheric processes and topographic features, which would be important for small-scale phenomena such as foehn.
One way to overcome these limitations is to combine AWS measurements and reanalysis data using statistical or machine learning techniques. A classification of foehn based on AWS data serves as the response (target/outcome/labels) for a supervised model that uses reanalysis data as explanatory variables (inputs/covariates). Once the relationship between the two sets of data has been learned, the models can be used to predict the expected state (‘foehn’ or ‘no foehn’) for periods for which reanalysis data is available but AWS measurements are not. This is also known as (statistical) downscaling or post-processing and has been used for foehn modeling in some variations: For nowcasting foehn at a station in Switzerland (Altdorf), Sprenger et al. (2017) use data from a local 7 km analysis data set (COSMO-7) – but only to get consistent inputs for their adaptive boosting algorithm (AdaBoost). Laffin et al. (2021) use both ERA5 and the Regional Atmospheric Climate Model 2 (RACMO2; Wessen and Laffin 2022) combined with tree-based gradient boosting (XGBoost; Chen and Guestrin 2016) to predict foehn on the Antarctic Peninsula. XGBoost is also used by Mony et al. (2021) to investigate future changes to foehn frequency in Switzerland.
This study proposes a novel probabilistic approach that combines unsupervised and supervised machine learning methods to bridge the gap between in-situ Automatic Weather Station (AWS) measurements and ERA5 reanalysis data to diagnose foehn. This combined approach allows us to reconstruct long-term, high-resolution foehn conditions over several decades. ERA5 data enables us to reconstruct the probability of foehn occurrence with hourly temporal resolution dating back to 1940, long before AWS were installed.
Our approach also allows to identify potential long-term changes in foehn occurrence in the European Alps from this high-resolution reconstruction. The validity of this combined approach is demonstrated by applying it to six stations located both north and south of the main Alpine ridge to show the method’s effectiveness for both north and south foehn wind situations.
Figure 1 shows a schematic representation of the proposed approach. For all six locations, data are available from an AWS at the target location as well as from a nearby mountain station for the last 14–22 years (Sec. 2.1) on a 10 min temporal interval. A Gaussian mixture model (unsupervised learning; Sec. 3.1) is used for foehn classification. The result is then aggregated into binary time series (‘foehn’/‘no foehn’) with an hourly temporal resolution to match the resolution of the ERA5 data used in the next step. After combining the different data sources, supervised learning (Sec. 3.2) is used to find the relationship between a variety of interpolated and derived variables from ERA5 (Sec. 2.2) and the classified events. Once these statistical models have been estimated, foehn can be reconstructed (Sec. 3.3) for the whole period from 1940–2022, allowing to investigate possible trends and/or seasonal changes over the last decades (Sec. 3.4).
2 Data
Section 2.1 describes the measurement data utilized for foehn classification along with the study area and the target stations. Section 2.2 explains the reanalysis data set and its pre-processing for the supervised learning method, along with the reconstruction process.
2.1 In-situ measurements
This study utilizes data from six AWSs situated across Switzerland and the western part of Austria, all positioned at the bottom of valleys known to be affected by foehn winds. Four of these stations are located north of the main Alpine ridge, while two are located in the canton of Ticino (Switzerland) south of the main Alpine ridge. Whilst the stations north of the Alps are prone to south foehn, the stations south of the Alps are known for the presence of north foehn.
An additional AWS upstream near the crest of the main Alpine range improves the accuracy of the foehn classification (Plavcan et al., 2014). The stations Innsbruck and Ellbögen utilize data from Sattelberg, the remaining four stations in Switzerland use observations from station Gütsch.
All stations provide data on mean wind speed, wind direction, air temperature, and relative humidity at a 10 min temporal resolution. Table 1 displays the locations and data availability of these stations, while Figure 2 depicts a map illustrating their geographical position and the surrounding topography.
| Type | Location | Data availability | |
| Gütsch (Andermatt)1 | crest | 46.653N/8.616E 2286m | 2005-01-01–2023-12-31 (95.3%) |
| Altdorf1 | south | 46.890N/8.620E 438m | 2005-01-01–2022-12-31 (78.1%) |
| Montana1 | south | 46.290N/7.460E 1423m | 2005-01-01–2022-12-31 (77.1%) |
| Comprovasco1 | north | 46.460N/8.935E 576m | 2005-01-01–2022-12-31 (85.8%) |
| Lugano1 | north | 46.004N/8.960E 205m | 2005-01-01–2022-12-31 (90.2%) |
| Sattelberg2 | crest | 47.011N/11.479E 2107m | 2006-01-01–2022-12-31 (75.0%) |
| Ellbögen2 | south | 47.200N/11.430E 1080m | 2006-01-01–2022-12-31 (92.0%) |
| (Universität) Innsbruck3 | south | 47.260N/11.385E 578m | 2009-06-21–2022-12-30 (99.1%) |
2.2 ERA5 reanalysis
This study makes use of ERA5 reanalysis data, which is publicly accessible via the Copernicus climate data store (Hersbach et al., 2023a, b). ERA5 offers four-dimensional gridded data with an hourly temporal resolution (starting from 1940) on a spatial grid of ( for Central Europe).
The data from 90 different fields (30 single level fields and 60 pressure level fields; see App. A) are bilinearly interpolated to the geographical location of the six stations of interest. Based on the 90 interpolated values, a series of derived variables is calculated, including vertical temperature gradients and level thickness, resulting in a total of 155 variables. These 155 variables are referred to as the ‘direct’ variable set as they solely rely on information retrieved directly from the geographical location of the corresponding target station.
Since foehn is driven by large-scale synoptic processes, it might be insufficient to rely only on data at the target location. Therefore, additional information from the surrounding is incorporated by extracting data from ERA5 at a series of neighboring points in a ‘star’ formation around the target location. In combination with the direct variables, a list of additional derived variables is calculated, such as spatial temperature differences and spatio-temporal pressure differences, among many others. Finally, the first and second order harmonics of the day of the year are included to capture seasonal variation. In total, this yields 497 variables: four harmonics, 155 direct variables, 136 spatial variables, 120 temporal variables and 82 spatio-temporal variables (e.g., temporal changes of spatial differences). This expanded set is referred to as the ‘full’ variable set.
Figure 2 shows the target locations (C; center) and
their neighboring points used. These neighboring points are positioned
geographically relative to the target station upstream (U) and
downstream (D) of the main foehn wind direction as well as to the right
(R) and left (L) of it. While the interpolated information from
the target location itself (C) is always used as possible covariate for
the statistical models, the values interpolated at the neighboring points are
not directly employed but are instead used for the calculation of the
derived/augmented variables. Further details regarding the
construction of the neighboring ‘star’ can be found in
Appendix A.
3 Methodology
Section 3.1 introduces the unsupervised learning model used for foehn classification, followed by the supervised learning models in Section 3.2. The results from Section 3.2 are then used to reconstruct hourly foehn occurrence over the past decades (Sec. 3.3) which is analyzed employing season-trend decomposition in Section 3.4.
3.1 Unsupervised learning: Mixture model for foehn classification
As direct foehn measurement do not exist, the data from the AWSs are currently unlabeled. Therefore, a mixture model is employed to distinguish between the ‘foehn’ and ‘no foehn’ events:
denotes the posterior
probability for a ‘foehn’ observation at a specific time and station, which is
modeled as a function () of the 10 min in-situ measurements
(Sec. 2.1).
We employ a two-component Gaussian mixture model (Grün and Leisch, 2008)
with concomitants, closely following the method proposed by Plavcan et al. (2014)
implemented in the R package foehnix (Stauffer, 2023).
The prerequisite condition for an observation to be used for classification is that the wind direction at the location of interest falls within the prevailing foehn direction at the target location, and that wind from a specific direction is also prevalent at the corresponding crest station at the same time (details in Tab. 4). Only the periods matching this precondition are used for estimating the Gaussian mixture model, while is set for all remaining observations.
The underlying concept involves that two unobservable Gaussian components (or clusters) exist, one describing ‘foehn’ conditions and the other ‘no foehn’ conditions. To distinguish between these two components, a main covariate is required. In this study, the potential temperature difference () between the valley station () and the crest station () is used, employing a simple dry adiabatic temperature reduction based on the difference in altitude (; cf.Tab. 1 & 4). During foehn events, the air descends on the leeward side of the mountains, resulting in a well-mixed atmosphere with neutral stratification, hence a potential temperature difference close to zero.
However, the potential temperature difference alone might not be sufficient to adequately separate the two states. Therefore, an additional concomitant model is employed to weigh the two components conditional on additional covariates. In this study, binary logistic regression is used for the concomitant model, employing relative humidity and mean wind speed as additional covariates. Models with this specification have been shown to work well for stations in the Alpine region (e.g., Plavcan et al. 2014; Plavcan and Mayr 2015). Figure 3 provides an illustration of this model, depicting the use of to separate the two components (‘foehn’/‘no foehn’) and the effect of the concomitant model on the joint density. More details can be found in Appendix B.
Once estimated, the mixture model provides the posterior probability for each 10 min interval where the required in-situ measurements are available (Eqn. 1). As these foehn ‘observations’ need to be combined with ERA5 in the next step, the results are upscaled to an hourly temporal resolution (). Therefore, the following basic assumption is applied: Each hour is considered a ‘foehn’ event if at least half of the 10 min posterior probabilities within that hour (e.g., 00:10–01:00 UTC for 01:00 UTC) are , otherwise the hour is considered a ‘no foehn’ event. If fewer than 4 out of 6 individual 10 min probabilities are available, the hour is excluded. This is similar to Gutermann et al. (2012) and Mony et al. (2021) who employ a ‘four out of six rule’.
3.2 Supervised learning: Modeling foehn probability
The outcome of the previous section is a binary time series, with each hour labeled as either a ‘foehn’ or ‘no foehn’ observation (Eqn. 2). This serves as the response variable in the supervised machine learning model, using the ERA5 data (Sec. 2.2) as input variables. The goal is to capture the relationship between the two data sets with model of the form
where the probability of the binary response is modeled as a function of the available information extracted from ERA5. can be any learner suitable for a binary response such as logistic regression, decision trees, random forests, or neural networks, to mention a few. For this study, three different learners/models are employed:
In order to investigate the possible benefits of incorporating large-scale information from the stations’ neighborhood, two variations of each learner are considered: one utilizing the ‘full’ set of 497 variables and one only using the ‘direct’ set of 155 variables (Sec. 2.2).
To account for location and time of day, separate models are estimated for each of the six stations (Tab. 4) for each hour of the day ( UTC, UTC, …, UTC), resulting in a total of 864 models. Depending on the station, the training data for these models include 10–18 years of data (see Tab. 1). In addition, a six-fold cross-validation (CV) is performed using a fixed period of 12 years (2011–2022) where, in each fold, two consecutive years are left out as test data. Ellbögen and Innsbruck are missing one fold (with test data 2013–2014) where no measurements from the crest station are available and thus the classification is not possible.
3.3 Long-term foehn reconstruction
Once the models from the previous section are estimated, they can be applied to the entire ERA5 period available. Although this is a prediction from a statistical perspective, it is termed ‘reconstruction’ in this article, as these predictions are applied backwards in time. The result is an hourly time series of foehn probabilities from January 1, 1940 to December 31, 2022 (83 years).
This reconstruction can fill possible gaps in historic records during the lifetime of an AWS and extend ‘foehn observations’ to times prior to or beyond the installation period of the AWS. The results may also serve as valuable input for investigating the effects of foehn, e.g., in ecology where the warming and drying effects of frequent foehn events can significantly impact flora and fauna or favor the occurrence of forest fires.
Moreover, the reconstruction offers the possibility to analyze foehn occurrence from a climatological perspective: Did the occurrence of foehn increase/decrease along with the changing climate over the decades? Are there changes in the seasonal or diurnal patterns? These questions are investigated in more detail in the next section.
3.4 Season-trend decomposition
The comprehensive reconstructed data set allows to study foehn occurrence in a climatological context. For this analysis, the hourly probability (Eqn. 3) is aggregated by (i) taking the highest probability per day (0000 UTC–0000 UTC), before (ii) calculating monthly averages. The resulting time series contains “monthly means of the daily maxima”, which are then modeled using a season-trend decomposition.
Due to the nature of the data there is a large year-by-year but also within-year variability depending on the prevailing weather situation. To decompose the signal, a season-trend decomposition is employed separating the signal into long-term changes and a remainder component containing the residual variability.
In this study, the regression-based decomposition of Dokumentov and Hyndman (2022) is used, which also provides confidence intervals for the estimated season and trend components. The model for the monthly mean foehn probabilities assumes an additive decomposition into a smoothly changing long-term trend , a smoothly changing seasonal component , and a remainder :
where is the seasonal frequency (i.e., monthly) and is
the sample size (83 years 12 months = 996). The model is estimated
via the R package stR (Dokumentov and Hyndman, 2023).
4 Results
First, this section investigates which insights can be gained from the reconstruction about the foehn occurrence at the six different target stations. Different temporal scales are considered for this, namely: Inter-annual changes in the foehn probabilities in Section 4.1, long-term trends and seasonal patterns in Section 4.2, as well as changes in the diurnal patterns across decades in Section 4.3. All of these results are based on the reconstruction using the ‘lasso’ learner with the ‘full’ covariate set for the full time period (without cross-validation).
Second, the performance of the supervised learning model is assessed under different model specifications, namely: Using the ‘full’ set of all 497 variables vs. the 155 ‘direct’ variables only in Section 4.4 and comparing the performance of the three supervised learners (lasso, stability selection, extreme gradient boosting) in Section 4.5. All of these results are based on out-of-sample Brier scores obtained in a six-fold cross-validation.
4.1 Temporal changes: Average annual foehn probabilities
The primary outcome of this study is the complete reconstruction of hourly foehn probabilities over years (see Sections 3.1–3.3), yielding time series with observations. To carve out inter-annual variations in the foehn probabilities, we aggregate the reconstructed data by taking the daily maximum of (Eqn. 3) before calculating annual means. This can be interpreted as the average probability of observing a foehn event on any given day within that year.
Figure 4 contains the result for all six stations, with Ellbögen exhibiting the highest mean annual probabilities (on average ), while Altdorf and Innsbruck show the lowest (on average and respectively). Additionally, the annual mean of daily maxima from the classification is shown for years with at least 80% of measurements available at the AWSs. The results show an overall good agreement between the two signals from the reconstruction and from the classification, with some larger gaps due to data availability as well as some noticeable differences for specific stations in particular years.
The reconstruction reveals a pronounced inter-annual variability, with certain years exhibiting a much higher annual mean than the long-term average, whilst others distinctly fall below. This variability is not random, as one can see similar patterns among the four south-foehn stations. For instance, all four stations show unusually high mean probabilities for 1951 and 1972. Similarly, the two north-foehn stations exhibit a similar temporal behavior over time.
Moreover, the figure suggests a possible increase in the annual mean foehn probability for south-foehn stations between 1940 and 1980. Hence this question is investigated in more detail in the next section.
4.2 Temporal changes: Climatological trends and seasonal patterns
In this section, the analysis from the previous section is taken a step further. Rather than focusing on the inter-annual variation the goal is to bring out the long-term climatological trends and changes in the seasonal patterns. Hence the reconstructed hourly time series are again aggregated but to monthly (rather than annual) means of the daily maxima of . Based on the season-trend decomposition outlined in Section 3 Figure 5(a) illustrates the resulting smooth trends () along with the corresponding confidence intervals. Figure 5(b) depicts the corresponding smoothly varying seasonal signals () averaged over decades for visual purposes (where decade 1940 corresponds to the years 1940–1949 etc.).
The estimated trends (Fig. 5a) show a significant increase for the two stations in Western Austria (Innsbruck, Ellbögen) between 1940 and 1980, and a plateau thereafter. All other stations show a linear increase over the study period. For four of the six stations (all except Montana and Comprovasco) the smooth trend differs significantly from a constant.
Figure 5(b) shows the analysis of the seasonal changes which reveals the different characteristics between north-foehn stations and south-foehn stations. The two stations located south of the main Alpine ridge, Comprovasco and Lugano, show one distinct maximum in spring and a minimum during autumn. This pattern is stable over the entire study period and no changes in the seasonal pattern () are found. The picture looks different when focusing on the four south-foehn stations which all show two maxima in spring and autumn with lower probability of foehn occurrence during summer and winter. Although not significant, the season-trend decomposition indicates an increase in the probability of foehn in spring (April, Mai) as well as in autumn (October, November) with a slight decrease in late summer (August, September).
4.3 Temporal changes: Intra-daily patterns
For certain applications, information about intra-daily patterns and their changes over time can be of great interest. With the hourly temporal resolution of the reconstruction, such insights are now possible across several decades.
Figure 6 shows Hovmøller diagrams for Ellbögen, depicting the decadal mean probability per time of day and month. Despite pronounced variability between the decades, the plot supports the previous findings showing an overall increase over time with strongest increase in spring (April, May) and in autumn (October, November). In addition, this visualization gives insights into the diurnal pattern. Generally, foehn occurrence in Ellbögen is more likely during the day (around 1000 UTC–2200 UTC) in spring and autumn, the period where indication for a certain increase was found (Sec. 4.2). The minimum average foehn probability is in the early morning and tied to the length of the night. The time of the minimum varies somewhat interdecadally.
4.4 Model performance: Benefit of full covariate set
As described in Section 3.2, two variants of the supervised learning methods are estimated: one using only the 155 ‘direct’ variables and one using the 497 ‘full’ variables including large-scale atmospheric conditions (Sec. 2.2). For both variants, a six-fold CV is performed (Sec. 3.2).
To investigate the benefit of the ‘full’ covariate set, Figure 7 shows the Brier scores (BSs) for the ‘lasso’ model for all six stations. For each station and each variant, BSs are shown for the test data set (out-of-sample) as well as for the training data set (in-sample). This shows that the models based on the ‘full’ variable set clearly outperform the models based on the ‘direct’ variables only. Although the overall performance of the less complex models based on the ‘direct’ variable set is still decent, including the additional large-scale spatio-temporal information substantially improves the overall model performance.
In addition to the predictive skill, Figure 7 also shows the stability of the model with largest variance in the BSs visible on the test data sets in Ellbögen. On the training data sets the scores barely vary due to the large sample size (10 years, hourly data).
4.5 Model performance: Comparison of supervised learners
While the previous section focuses on the benefits of using more input data, this section compares the three different supervised learners described in Section 3.2. For simplicity, only the results for models based on the ‘full’ variable set are shown as they have been shown to outperform those only using the ‘direct’ variable set (Sec. 4.4).
Figure 8 illustrates that the BSs from the six-fold CV on the test set (out-of-sample) is comparable for all three supervised learners with only minor differences. The average BS is slightly lower for ‘lasso’ (), followed by ‘stabsel’ (), and ‘xgboost’ ().
On the training data (in-sample) the picture is similar for ‘lasso’ and ‘stabsel’ but ‘xgboost’ has much lower BSs. This indicates that ‘xgboost’ might be subject to some overfitting, despite careful tuning of the hyperparameters (App. C.3).
5 Discussion and outlook
Using a novel combination of unsupervised and supervised learning we are able to accurately reconstruct long-term foehn time series (starting from 1940) at hourly resolution. More specifically, foehn classification is first accomplished by (unsupervised) Gaussian finite mixture models based on AWS measurements and then linked to ERA5 data using binary supervised learners such as lasso, stability selection, or extreme gradient boosting. The resulting foehn reconstruction enables novel analyses, exemplified here by investigating long-term changes in trends, seasonal patterns, and diurnal cycles of foehn occurrence.
The season-trend decomposition based on the period 1940–2022 reveals that all six stations considered have either experienced a linear increase in foehn occurrence (probability) over the entire study period or an increase between 1940 and the early 1980s that leveled off afterwards. For four of the six stations it can be shown that this increase is significant. Although no significant change in the seasonality is found, the results for all south-foehn stations (Altdorf, Montana, Ellbögen, and Innsbruck) indicate a slight increase in the occurrence of foehn in spring and autumn, with a slight decrease in late summer.
The high quality of the foehn reconstruction is partially due to using a large set of predictor covariates that not only contains information at the target location but also includes additional large-scale atmospheric information from the stations’ surrounding. The benefits of this full set of covariates are similar for all three supervised learners considered: logistic regression with lasso regularization (‘lasso’), logistic-regression-based stability selection (‘stabsel’), and extreme gradient boosting (‘xgboost’). Lasso performs best in our application, closely followed by the other two learners. Some further improvements might be gained for xgboost if overfitting on the training data can be further reduced, e.g., by a different hyperparameter tuning strategy.
For a comparison to existing publications, we complement the Brier scores shown in the main paper by other popular scores for binary outcomes, such as the false negative rate (FNR, also known as miss rate), false positive rate (FPR, also known as false alarm rate), and percent correct (PC, also known as accuracy). Based on the best model (lasso with full covariate set), we obtain the following performances for Altdorf: // (FNR/FPR/PC). These align well with the existing literature and stand out for an exceptionally low FPR. Sprenger et al. (2017) report // and Mony et al. (2021) report //. Similarly for Lugano we obtain //, while Mony et al. (2021) report //. More details are included in the appendix (App. E).
The same holds for the foehn classification when compared to existing literature and the Swiss foehn index (SFI) operationally used at MeteoSwiss in terms of “average foehn hours per year”. The results from the Gaussian mixture model (Sec. 3.1) for Altdorf show an average of 482.4 h/year which aligns well with the SFI (458.4 h/year), as well as the results reported by Dürr (2008) (478 h/year), Jansing et al. (2022) (465.8 h/year), and MeteoSwiss (2024) (477 h/year). Montana exhibits 1007.4 h/year (SFI 904.7 h/year111Average based on the years 2009–2016 only.), while Lugano and Comprovasco show 644.7 h/year (SFI 563.0 h/year, MeteoSwiss (2024) 551 h/year) and 1077.4 h/year (SFI 953.4 h/year), respectively.
It should be emphasized that the excellent performance of the combined approach depends crucially on the high quality of the automatic foehn classification in the Gaussian finite mixture model. Thus, if the separation of the AWS measurements into ‘foehn’ or ‘no foehn’ components works well (as for the six stations presented), then different binary supervised learners are able to link this reliably to the ERA5 data. However, there may also be stations where a distinction into just two components might not be sufficient. For example, we have found this to be the case for Aigle, Switzerland, where a three-component mixture model appears to be necessary to separate light down-valley winds, strong humid (catabatic) outflows, and actual foehn situations (details not shown).
Computational details
[…]
Acknowledgments
The study is partly based on the preliminary work of Morgenstern (2020). The computational results presented here have been achieved (in part) using the LEO HPC infrastructure of the University of Innsbruck.
Appendix A ERA5 fields
As described in Section 2.2 information from 90 different ERA5 fields is bilinearely interpolated to (i) the target location of the different stations used (cf. 2.1) as well as to a series of neighboring points arranged in a ‘star’ formation to include large-scale information. The fields used are listed in Table 2 and 3.
| CDS ERA5 fields name | |
100m_u_component_of_wind |
100m_v_component_of_wind |
10m_u_component_of_wind |
10m_v_component_of_wind |
10m_wind_gust_since_previous_post_processing |
instantaneous_10m_wind_gust |
2m_dewpoint_temperature |
2m_temperature |
surface_pressure |
mean_sea_level_pressure |
eastward_gravity_wave_surface_stress |
northward_gravity_wave_surface_stress |
convective_precipitation |
large_scale_precipitation |
total_precipitation |
boundary_layer_dissipation |
boundary_layer_height |
charnock |
friction_velocity |
gravity_wave_dissipation |
high_cloud_cover |
low_cloud_cover |
medium_cloud_cover |
total_cloud_cover |
total_column_cloud_liquid_water |
surface_net_solar_radiation |
surface_net_thermal_radiation |
surface_sensible_heat_flux |
surface_solar_radiation_downwards |
surface_thermal_radiation_downwards |
| CDS ERA5 fields name | |
divergence |
geopotential |
potential_vorticity |
specific_humidity |
temperature |
u_component_of_wind |
v_component_of_wind |
vertical_velocity |
specific_cloud_liquid_water_content |
vorticity |
To keep the approach generic, these neighbouring points are ‘arbitrary’ points
not linked to stations or cities, but geometrically placed around the
target station relative to its location to the main mountain range.
Figure 2 (main article) shows these ‘stars’ which
are constructed as follows: First, the
orthogonal from the center location (station location; C) to the
main mountain range is computed. Along this orthogonal axis, two additional
points upwind of the foehn wind direction (U, UU) and two
points in the downwind direction (D, DD) are defined with a
radius of 1∘ and 2∘ from the center point. Second, four
points with a radius of 2∘ from the center point deviating by
45∘ from the orthogonal are added, defining points on the right hand
side (*R) and left hand side (*L) of the main foehn direction
in the upwind direction (UR, UL) and downwind direction
(DR, DL). The downwind direction for stations located north
of the main alpine range is towards north (south-foehn stations), for those
located south of the main alpine range the downwind direction is toward
south (north-foehn stations).
Based on these points, the ‘direct’ variable set (155 variables) as well as the ‘full’ variable set (497 variables) are calculated as described in Section 2.2 (main article). The extended (‘full’) variable set contains a series of spatial and spatio-temporal information such as:
surface pressure differences between
CtoU,UtoC,ULtoDRand others, -
potential temperature differences between
CtoU,UtoC,ULtoDRand others, -
differences in vertical temperature differences between
CtoU,UtoC,ULtoDRand others, -
sum of precipitation on the upwind side (sum of
U,UU,UR,UL) as well as on the downwind side (sum ofD,DD,DR,DL) as well and the difference between these two sums, -
mean cloud cover on the upwind side (sum of
U,UU,UR,UL) as well as on the downwind side (sum ofD,DD,DR,DL) as well and the difference between these two means.
Moreover, temporal variables are calculated including e.g., changes in the geopotential height, temperature, or changes in humidity on different levels over the past 3 hours or over the next 3 hours and variations thereof.
Appendix B Unsupervised learning: Two-component mixture model
For all eight AWSs used (Tab. 1; main article) observations are available for the most recent 10–18 years on a 10 min temporal resolution. Since foehn winds show a characteristic wind direction, only observations meeting a specific precondition are used for classification as described in Section 3.1. Table 4 shows the wind sectors used for each target site, as well as the amount of data used for classification.
| Stations, conditional wind direction, data used | |
| Altdorf |
Valley station: Altdorf (60∘–240∘)
Crest station: Gütsch, Andermatt (105∘–285∘) 20.0 % within sector, 57.1 % outside sector, 22.9 % removed |
| Innsbruck |
Valley station: Universität Innsbruck (130∘–230∘)
Crest station: Sattelberg (90∘–270∘) 5.8 % within sector, 49.3 % outside sector, 44.9 % removed |
| Ellbögen |
Valley station: Ellbögen (45∘–225∘)
Crest station: Sattelberg (90∘–270∘) 22.6 % within sector, 49.0 % outside sector, 28.4 % removed |
| Montana |
Valley station: Montana (30∘–100∘)
Crest station: Gütsch, Andermatt (120∘–180∘) 23.2 % within sector, 52.9 % outside sector, 23.9 % removed |
| Comprovasco |
Valley station: Comprovasco (270∘–30∘)
Crest station: Gütsch, Andermatt (290∘–89∘) 33.5 % within sector, 51.2 % outside sector, 15.4 % removed |
| Lugano |
Valley station: Lugano (270∘–80∘)
Crest station: Gütsch, Andermatt (290∘–89∘) 26.9 % within sector, 62.1 % outside sector, 10.9 % removed |
The (conditional) two-component Gaussian mixture model with concomitant variables Grün and Leisch (2008); Plavcan et al. (2014); Stauffer (2023) is defined as follows:
where the joint (mixed) density is the sum of two Gaussian components times the probability that an observation falls into the second component. The potential temperature is used to separate the two components with representing the distribution parameters of . The concomitant model modeling is a logistic regression model employing an intercept as well as relative humidity and wind speed () and can be written as:
Once the required parameters , are estimated, the final a-posteriori probability can be calculated using
Appendix C Supervised learning: Reconstruction
The hourly labeled data (Sec. 3.1, App. B) are are combined with the hourly data derived from ERA5 (Sec. 2.2, App. A), serving the training data for the supervised learning.
C.1 lasso: Logistic regression with lasso (L1) regularization
Due to the large number of possible covariates from ERA5 and the fact that many of these variables are highly correlated, penalization is required. This study employs lasso (least absolute shrinkage and selection operator) with L1 regularization, optimizing the following penalized log-likelihood:
with , where is the number of hourly foehn probabilities (Sec. 3.1) and the number of available covariates (Sec. 2.2).
To find the optimal tuning parameter , a 30-fold cross-validation
is performed which yields the optimal (regularized) set of regression coefficients
using the R package glmnet
[…]
C.2 stabsel: Logistic regression with lasso-based stability selection
As the lasso penalty shrinks more and more parameters to with increasing , this model can also be used for stability selection (Meinshausen and Bühlmann, 2010). The penalized regression model is estimated times for a series of values . The first parameters which enter the model () will be memorized.
In this study, this is repeated times, estimating the model on a random subset with only 50% of the available data. As the number of ‘foehn’ events can be rather low (0.6%–29% depending on station/time of day) a stratified random subset (bagging) is drawn each time containing the same proportion of ‘foehn’ and ‘no foehn’ events as the full data set. In each of the iterations, the name of the first covariates entering the model is stored. Those covariates which have been selected more than of the time (at least 121/200 times) are then used to estimate an additional unregularized logistic regression model.
1.
Whilst iteration is :
Draw stratified subsample of size
Estimate regularized coefficients for different s
Extract and keep the name of the first covariates entering the model
2.
Select the covariates which a selection frequency
3.
Estimate final (unregularized) logistic regression model with the covariates from Step 2
Stability selection allows to automatically identify the most important covariates among the ones available and to strongly reduce the complexity of the final model (Step 3) without losing much predictive performance.
C.3 xgboost: Extreme gradient boosting
Extreme gradient boosting (XGBoost; Chen and Guestrin 2016) is a supervised machine learning algorithm using a series of decision trees to best predict the labeled data (response) using a series of explanatory variables (covariates). Similar to random forests, XGBoost uses estimates and combines a series of decision trees (a weak learner). The main difference is that XGBoost uses a gradient decent method (thus the name) by optimizing the model based on the outcome of the previous tree by finding additional split points to improve the objective function.
Overly simplified, XGBoost starts with estimating one decision tree and evaluates its residuals (i.e., the yet unexplained part/misclassified data). These residuals are then used as input for the next iteration to estimate a new tree to improve the misclassification from the previous step.
A series of tuning parameters exist such as the number of iterations to be
performed, the maximum depth of an individual tree, or the learning rate
to update the parameters in each iteration. Choosing these tuning parameters
wrong can lead to overly conservative models (not flexible enough) or models
which overfit the data, both of which can result in a poor (out-of-sample)
predictive performance. One way to avoid these scenarios is to perform
proper cross-validation. We performed a (limited) grid search using
the R package xgboost (Chen et al., 2023) with
the following parameter space:
eta, learning rate: -
max_depth, maximum depth of each weak learner (tree): -
min_child_weight, minimum sum of instance weight needed in a child: -
gamma, minimum loss reduction required:
This yields an overall number of unique combinations. For each combination, a 10-fold cross-validation using random stratified subsamples and a maximum of boosting iterations is performed. It is then evaluated which parameter set and boosting iteration resulting in the lowest average out-of-sample error to estimate one final model using all data. A summary of the selected tuning parameters is shown in Table 5. All models are estimated using the ‘gbtree’ booster and the binomial logistic objective function.
| mean | median | range | |
| subsample | 0.500 | 0.5 | 0.5-0.5 |
| eta | 0.200 | 0.2 | 0.2-0.2 |
| max_depth | 11.602 | 10 | 5-20 |
| min_child_weight | 3.423 | 4 | 2-6 |
| gamma | 2.868 | 2 | 2-10 |
| nround | 19.997 | 20 | 19-20 |
Appendix D Long-term reconstruction
Figure 4 (main article, Sec 4.1) only shows results of the reconstruction based on the ‘lasso’ model with the ‘full’ variable set. Figure 9 shows the same information for all six stations including the results of all three models (lasso, stabsel, xgboost) and both variants using either the ‘direct’ variable set or ‘full’ variable set.
The results of these six versions (three models, two variants each) show differences but agree in large parts, showing similar features across the years.
Appendix E Model performance/model comparison
Analogous to Figures 7 and 8 (main article; Sec,4.4 & 4.5) Figure 10 shows average Brier scores from six-fold CV for all three models and model variants (using the ‘direct’ or ‘full’ variable set). As discussed in the main article, the ‘xgboost’ model might be subject to some overfitting (training) as it always shows lowest scores across all stations and both model variants.
In addition, event-based metrics are used in the discussion (Sec 5) to compare the results of this study to existing literature. Table 6 contains the results for all six stations and both model variants (‘direct’/‘full’ variable set) for all three supervised methods (‘lasso’/‘stabsel’/‘xgboost’).
| Station | Variable set | FNR | FPR | PC |
| Altdorf | direct | // | // | // |
| full | // | // | // | |
| Innsbruck | direct | // | // | // |
| full | // | // | // | |
| Ellbögen | direct | // | // | // |
| full | // | // | // | |
| Montana | direct | // | // | // |
| full | // | // | // | |
| Comprovasco | direct | // | // | // |
| full | // | // | // | |
| Lugano | direct | // | // | // |
| full | // | // | // |